find the slope 8y-2x=24
Answer by nerdybill(7384)

Arrange your equation into the "slope-intercept" form:
y = mx + b
where
m is the slope
b is the y-intercept
.
8y-2x=24
Adding 2x to both sides:
8y= 2x+24
Dividing both sides by 8:
y = (2/8)x + (24/8)
y = (1/4)x + 3
.
Therefore, the slope is 1/4